tax rate was 30% and dividend eangs were taxed at 12 mained in dividend payments i the coora .5%. O A. $106,200 O B $133,000 O C. $150 800 O D. $106.,400
Answer #1

Answer

--The correct answer is Option 'B' $ 133,000

A Earnings per share before taxes $                     3.60
B Tax rate 30%
C = A x B Income tax expense per share $                     1.08
D = A - C Earnings after tax $                     2.52
E Amount retained $                     1.00
F = D - E Amount distributed as dividend per share $                     1.52
G No. of shares                  100,000
H = F x G Total Dividend received by holder $        152,000.00
I Dividend earnings tax rate 12.50%
J = H x I Tax on dividend earnings $          19,000.00
K = H - J Value of dividend earnings after tax $        133,000.00
Correct Answer Option 'B': $ 133,000